OpenStudy (anonymous):

integrate 12sin^23x help- calc exam tomorrow

OpenStudy (anonymous):

sorry integrate 12sin^2 *3x

OpenStudy (anonymous):

\[12\sin^2(3x)\]?

OpenStudy (anonymous):

Remeber you half angle identities?

OpenStudy (dan815):

use chain rule

OpenStudy (anonymous):

\[\int\limits_{?}^{?}12\sin^2(3x)dx=\int\limits_{?}^{?}12(\frac{ 1 }{ 2 })(1-\cos(6x)dx\]

OpenStudy (dan815):

oh integrate not derivative lol

OpenStudy (anonymous):

answer is \[6x-\sin(6x)+c\]

OpenStudy (dan815):